# Enhancing Memory Efficiency in Large Language Model Training Through Chronos-aware Pipeline Parallelism

### v1

@@ 4.1. Chronos-Pipe: Chronos-aware Pipeline
A further theoretical analysis of Chronos-Pipe’s impact on activation storage requirements and P2P communication is provided. For simplicity, we assume $T_{bwd}=2T_{fwd}$ and set $T_{c}$ for a single P2P communication to zero unless stated otherwise. With these assumptions, basic unit of time $T_{unit}$ in the pipeline scheduling diagram (e.g., Fig.  5(a)) is $\frac{T_{fwd}}{2p}$, and the activation storage for each task is $\frac{m_{a}}{2p}$.
@@
In Chronos-Pipe, activation storage can be estimated by analyzing the activation lifespan, approximately 75% of $m_{a}$. As shown in Fig.  5(a), Stage 0 launches chunk 2 of the fourth micro-batch before releasing the activation storage from chunk 2 of the third micro-batch. Since tasks on a single stage execute in cycles of $6T_{unit}$, the accumulated activation storage for chunk 2 can be represented as $\lceil\frac{T_{life\_chunk2}}{6T_{unit}}\rceil$, where $T_{life\_chunk2}$ denotes the lifespan of chunk 2’s activations, precisely the time from the completion of its forward pass to the beginning of the corresponding backward pass at PP stage 0. Similarly, the accumulation of activation storage for chunk 1 can be expressed as $\lceil\frac{T_{life\_chunk1}}{6T_{unit}}\rceil$. Additionally, we observe that the forward execution times for both chunks $T_{fwd\_chunk1}$, $T_{fwd\_chunk2}$ are each $pT_{unit}$, while the backward execution times $T_{bwd\_chunk1}$, $T_{bwd\_chunk2}$ are each $2pT_{unit}$. There is also an extra interval $T_{fwd\_interval}$ between the forward chunks, derived from steps 1, 2, and 3 in Fig.  5(a) as $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$. Similarly, an additional interval $T_{bwd\_interval}$ exists between the backward chunks, derived from steps 4, 5, and 6 in Fig.  5(a) as $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$. Thus, for Stage 0, the lifespans of the activations for chunk 1 and chunk 2, denoted $T_{life\_chunk1}$,$T_{life\_chunk2}$ can be expressed as:
@@
As a result, the peak activation storage requirements for chunks 1 and 2 are $\lceil\frac{2}{3}+\lceil\frac{(p-3)}{6}\rceil+\lceil\frac{(2p-3)}{6}\rceil+\frac{p}{2}\rceil\frac{m_{a}}{2p}$,$\lceil\frac{(3p-2)}{6}\rceil\frac{m_{a}}{2p}$ respectively. When P is large, these values approach ${}\frac{m_{a}}{2},\frac{m_{a}}{4}$, making the total peak activation storage approximately 75% of $m_{a}$.
In Chronos-Pipe, when the task is divided into v chunks, the bubble overhead from P2P communication is approximately v times that of the 1F1B schedule. The scheduling of Chronos-Pipe with two chunks under the condition $T_{c}\neq 0$ is considered to illustrate this. As shown in Fig.  5(b), the Chronos-Pipe execution time can be divided into three phases: $T_{warmup},T_{steady}$ and $T_{cooldown}$.
@@
$\bullet$ The warmup phase primarily includes the forward pass of the first micro-batch, calculated as $(2p+2)T_{unit}+(2p+1)T_{c}$.
@@
$\bullet$ The Steady phase covers the forward pass of chunk 1 for the remaining $(m-2)$ microbatches, totaling $(m-2)$$(6T_{unit}+8T_{c})$.

### v2

@@ Appendix A Theoretical Analysis of T-Pipe
A further theoretical analysis of T-Pipe on activation storage and bubble ratio is provided as follow. For simplicity, we assume $T_{bwd}=2T_{fwd}$, basic unit of time $T_{unit}$ in the pipeline scheduling diagram (e.g., Fig.  16) is $\frac{T_{fwd}}{2p}$, and the activation storage for each task is $\frac{m_{a}}{2p}$.
@@
In T-Pipe, activation storage can be estimated by analyzing the activation lifespan, approximately 75% of $m_{a}$. As shown in Fig.  16, Stage 0 launches chunk 2 of the fourth micro-batch before releasing the activation storage from chunk 2 of the third micro-batch. Since tasks on a single stage execute in cycles of $6T_{unit}$, the accumulated activation storage for chunk 2 can be represented as $\lceil\frac{T_{life\_chunk2}}{6T_{unit}}\rceil$, where $T_{life\_chunk2}$ denotes the lifespan of chunk 2’s activations, precisely the time from the completion of its forward pass to the beginning of the corresponding backward pass at PP stage 0. Similarly, the accumulation of activation storage for chunk 1 can be expressed as $\lceil\frac{T_{life\_chunk1}}{6T_{unit}}\rceil$. Additionally, we observe that the forward execution times for both chunks $T_{fwd\_chunk1}$, $T_{fwd\_chunk2}$ are each $pT_{unit}$, while the backward execution times $T_{bwd\_chunk1}$, $T_{bwd\_chunk2}$ are each $2pT_{unit}$. There is also an extra interval $T_{fwd\_interval}$ between the forward chunks, derived from steps 1, 2, and 3 in Fig.  16 as $(3+6\lceil\frac{(p-3)}{6}\rceil-p)T_{unit}$. Similarly, an additional interval $T_{bwd\_interval}$ exists between the backward chunks, derived from steps 4, 5, and 6 in Fig.  16 as $(3+6\lceil\frac{(2p-3)}{6}\rceil-2p)T_{unit}$. Thus, for Stage 0, the lifespans of the activations for chunk 1 and chunk 2, denoted $T_{life\_chunk1}$$T_{life\_chunk2}$ can be expressed as:
@@
As a result, the peak activation storage requirements for chunks 1 and 2 are $\lceil\frac{2}{3}+\lceil\frac{(p-3)}{6}\rceil+\lceil\frac{(2p-3)}{6}\rceil+\frac{p}{2}\rceil\frac{m_{a}}{2p}$$\lceil\frac{(3p-2)}{6}\rceil\frac{m_{a}}{2p}$ respectively. When P is large, these values approach $\frac{m_{a}}{2},\frac{m_{a}}{4}$, making the total peak activation storage approximately 75% of $m_{a}$.
In addition, T-Pipe exhibits a bubble ratio similar to 1F1B. As depicted in Fig.  16, its execution comprises three phases: $T_{warmup},T_{steady}$ and $T_{cooldown}$.
@@
$\bullet$ The warmup phase includes the forward pass of the first micro-batch, calculated as $2pT_{unit}$.
@@
$\bullet$ The Steady phase covers the forward pass of chunk 1 for the remaining (m1)(m-1) microbatches, totaling $6(m-1)T_{unit}$.